# The generalized Ramsey number $f(n,5,8)=\frac{6}{7}n+o(n)$

### v1

@@ 1. Introduction
Gomez-Leos, Heath, Parker, Schweider and Zerbib [17] recently proved the lower bound $f(n,5,8)\geq\frac{6}{7}(n-1)$. In this paper we give an alternate proof for this lower bound, since we believe it illuminates what must be done to get a matching upper bound. But our main contribution is to obtain a $(5,8)$-coloring using $\frac{6}{7}n+o(n)$ colors. To do this we will actually separately apply certain results of Delcourt and Postle [9] as well as Glock, Joos, Kim, Kühn and Lichev [14]. The results in [9] and [14] are quite similar in the big picture, but there are important technical differences that give each one an advantage over the other in certain instances.
@@ 2. Gadget packing and conflict-free hypergraph matchings
Very recently, Glock, Joos, Kim, Kühn, Lichev and Pikhurko [15] proved Conjecture 1 for the case $\ell=6$ by finding the limit. The proof is significantly more complicated but bears a resemblance to the sketch we provided above for Glock’s proof [13] of the case $\ell=5$. In particular, the lower bound in [15] is obtained by packing an appropriately defined gadget, i.e. a hypergraph which has some sharable pairs. The upper bound follows from establishing a few constraints like (2) and (3) linear programming (and the optimal solution to the linear program tells us what gadget to pack). However, to avoid all possible $(6,4)$-configurations, one must avoid certain configurations consisting of several gadgets intersecting in certain ways. For this, a result like Pippinger and Spencer [22] is not sufficient, and so they use conflict-free hypergraph matchings instead.
@@ 3. Lower bound
Consider a $(5,8)$-coloring of $K_{n}$ using color set $C$. We will use this coloring to define a certain partition of the edges $E(K_{n})$. We start with the partition $\mathcal{P}^{\prime}$ into monochromatic components. In a $(5,8)$ coloring a monochromatic component has at most three edges. For the rest of this proof whenever we say “component” we mean a monochromatic component. We assume there are no monochromatic triangles so each part of $\mathcal{P}^{\prime}$ is either an isolated edge, 2-path, 3-path, or 3-star. For each part $P\in\mathcal{P}^{\prime}$ consisting of a component in color $c$, we say $P$ hits each pair $(v,c)$ where $v$ is a vertex in the component. Clearly, each pair $(v,c)$ is hit by at most one part.
@@
Now we start to describe the merging. First, for any part $P=\{xy,xv,xy\}\in\mathcal{P}^{\prime}$, i.e. a 3-star component, we see that $\mathcal{P}^{\prime}$ also has singleton parts $\{uv\},\{uy\},\{vy\}$ (if any of these edges were in a larger component we would have a $(5,8)$-violation). See Figure 2(1). We merge $P$ with these singletons forming a new part $Q$ with six edges. Note that two distinct 3-stars can share at most one vertex since otherwise we have a $(5,8)$-violation. Thus each singleton part in $\mathcal{P}^{\prime}$ gets merged with at most one 3-star. In addition to all the 10 pairs hit by the parts we merged to form $Q$, we also say that $Q$ hits $(x,\text{green}),(x,\text{red}),(x,\text{orange})$ (see Figure 2(1)). Any scenario in which a pair is hit twice implies a $(5,8)$-violation. We merge all 3-stars with their respective singletons to form a new partition. We abuse notation and call the new partition $\mathcal{P}^{\prime}$, and we will keep doing so after even more merging until we reach our final partition $\mathcal{P}$.
Next, for any part $P=\{ux,xy,yv\}\in\mathcal{P}^{\prime}$, i.e. a 3-path component, we see that $\mathcal{P}^{\prime}$ also has singleton parts $\{uy\},\{xv\}$. See Figure i.e. 2(2). We merge $P$ with these singletons forming a new part $Q$ with five edges. We say the new part hits the additional pairs $(x,\text{green}),(y,\text{red})$. Next, for any pair of parts $P=\{xu,uv\},P^{\prime}=\{xy,yv\}\in\mathcal{P}^{\prime}$, i.e. a pair of 2-paths sharing endpoints, we will merge $P,P^{\prime}$ and the singleton part $\{xv\}$. See Figure i.e. 2(3). We say the new part hits the additional pairs $(u,\text{green}),(y,\text{green})$. Next, for any pair of parts $P=\{xy,yu\},P^{\prime}=\{xu,uv\}\in\mathcal{P}^{\prime}$ we will merge $P,P^{\prime}$ and the singleton part $\{xv\}$. See Figure i.e. 2(4). We say the new part hits the additional pair $(u,\text{green})$.
@@
For $i=1,\ldots,8$ let $x_{i}$ be the number of parts in $P$ that look like Figure 2$(i)$. Then summing the edges we have
@@
Now since each pair $(v,c)$ is hit by at most one part of $\mathcal{P}$ we have
@@
$=\frac{19}{7}x_{1}+\frac{10}{7}x_{2}+\frac{10}{7}x_{3}+\frac{3}{7}x_{4}+\frac{6}{7}x_{5}+\frac{1}{7}x_{6}+\frac{2}{7}x_{8}$
$\geq 0.$ (5)
@@ 4.1. Intuition
We will need to take care how the gadgets in our packing intersect. We need almost all the edges of $K_{n}$ to be covered by edge-disjoint gadgets, and if we are careless we might end up with two gadgets intersecting as shown in Figure 3(2). In that figure the red and blue edges come from one gadget, and the green and orange from a second gadget. If the two gadgets intersect as shown we see a $(5,8)$-violation. To avoid this kind of complication we will only allow a pair of gadgets to share a pair of vertices if that pair is a nonedge in both gadgets, as shown in Figure 3(3). In order to ensure we can still cover almost all edges of $K_{n}$ with our gadgets, we will need to make sure that there are not too many of these pairs of vertices that get used as nonedges of gadgets (such pairs will never be covered by our packing). Thus, before we try to get our packing, we will designate a small set of edges to be our shareable edges (see Figure 3(1)). When we pack a gadget, we will make sure the non-edges of the gadget are all shareable, and the edges of the gadget are not shareable. This way we can allow two gadgets to share two vertices as long as this pair of vertices is a shareable edge and therefore a non-edge in each of the gadgets. We will call this gadget in Figure 3(1) $\Gamma$. We will define $\Gamma$ more formally later, but it will be a set of edges colored according to the pattern shown in Figure 3(1) but with the colors possibly permuted, with the dotted edges always being sharable. The formal definition of $\Gamma$ will also involve additional rules about how the colors are chosen, which we start to describe next.
@@ 4.3. Overview of the coloring
Our construction will start by identifying our set of shareable edges, and non-hitting colors for each vertex. Then we would like to color almost all the edges (excluding shareable edges and a few other edges) by packing edge-disjoint gadgets. These gadgets must be packed to obey the rules for shareable edges and the colors chosen to obey the rules for non-hitting colors. We will use two disjoint sets of colors $C_{A1}$ and $C_{A2}$ for these gadgets, and the colors in $C_{A1}$ and $C_{A2}$ will play different roles. For $c\in C_{A1}$, $c$ will only be allowed to play the role of blue in Figure i.e. 2(7), and so the color class for $c$ will consist of vertex-disjoint 1-paths (i.e. a matching in $K_{n}$). For $c\in C_{A2}$, $c$ will only be allowed to play the role of red or green in Figure i.e. 2(7), so the color class for $c$ will consist of vertex-disjoint 2-paths. In fact, to get our asymptotically optimal bound on the number of colors we will need each color in $C_{A1}$ to induce an almost-perfect matching and each color in $C_{A2}$ to induce a set of 2-paths covering almost all vertices. When we pack the gadgets the colors will also be chosen to obey certain other rules to avoid $(5,8)$-violations. This part of our construction will be called Phase A, and the proof that it works will be an application of Theorem 2. In particular we will define an auxiliary hypergraph $\mathcal{H}_{A}$ and conflict system $\mathcal{C}_{A}$ such that a $\mathcal{C}_{A}$-free matching in $\mathcal{H}_{A}$ corresponds to a partial $(5,8)$-coloring. The advantage of using Theorem 2 here is that this theorem has some extensions which we will use (see Section A). These extensions allow us to prove that the coloring obtained in Phase A has some nice quasirandom properties which will allow us to do the next phase. In Phase B we will color all edges that remain uncolored after Phase A. This will be done by applying Theorem 3 to a new hypergraph $\mathcal{H}_{B}$ and conflict system $\mathcal{C}_{B}$.
@@ 5.1. Setting up our application of Theorem 2
We will apply Theorem 2 with $\varepsilon=\varepsilon_{A}$, and at certain points of our proof we will see inequalities that $\delta,\varepsilon_{A}$ must satisfy. We will say $\delta,\varepsilon_{A}$ are chosen chosen such that
@@
The conflict system $\mathcal{C}_{A}$ for $\mathcal{H}_{A}$ of course has vertex set $E(\mathcal{H}_{A})$, i.e. the set of gadgets. The conflicts (edges) of $\mathcal{C}_{A}$ will be certain sets of gadgets which together would create a $(5,8)$-violation or an alternating 4-cycle (a 4-cycle that is alternately colored using 2 colors).
@@
  • Type 5 conflicts (see Figure 5(5)) correspond to collections of 5 gadgets $\Gamma_{1},\ldots,\Gamma_{5}$ making a $(5,8)$-violation $S$ in the following way. $\Gamma_{1}$ has a triangle in $S$. $\Gamma_{2}$ and $\Gamma_{3}$ each have $c$-colored edge (blue) in $S$, and $\Gamma_{4},\Gamma_{5}$ each have a $c^{\prime}$-colored edge (green) in $S$. The $c$-colored edges are disjoint, as are the $c^{\prime}$-colored edges, and there is no alternating 4-cycle in $S$. In particular the $c$- and $c^{\prime}$-colored edges together make an alternately colored 5-path $(s_{1},s_{2},\ldots,s_{5})$ where $s_{1},s_{3},s_{5}$ are all on the triangle from $\Gamma_{1}$. This path could be oriented in a few ways with respect to the orientation of the triangle from $\Gamma_{1}$, and the figure shows one possible orientation.
@@
Indeed, by the previous observation $\Gamma$ cannot have a fourth vertex in $S$. For the 3 vertices $\Gamma$ has in $S$, both $c,c^{\prime}$ are either non-touching for these vertices or else they are already touching due to $\Gamma$. Thus $S$ only has one edge which could be colored either $c$ or $c^{\prime}$.
@@
In light of Proposition 1 and the fact that $C_{A}$ has $\frac{6}{7}n+O(n^{1-\delta})$ colors, our proof of Lemma 5 will be done if we show there is a large enough $\mathcal{C}_{A}$-free matching in $\mathcal{H}_{A}$. To do that, we start checking the conditions of Theorem 2.
@@ 5.2. Checking conditions of Theorem 2
We start verifying the claim with $j=4$. We will check one case in more detail and then provide a general argument for the rest. We will check Type 1 conflicts in detail, i.e. we will check that each gadget is contained in $O(d^{3})$ conflicts of Type 1. Fix a gadget $\Gamma_{1}$. A Type 1 conflict in $\mathcal{C}_{A}^{(4)}$ containing $\Gamma_{1}$ contains 3 other gadgets $\Gamma_{2},\Gamma_{3},\Gamma_{4}$ (see Figure 8(1)). Fixing $\Gamma_{1}$ and assuming we choose all the information for $\Gamma_{2},\Gamma_{3},\Gamma_{4}$ in that order, the choices we must make are summarized in Table 1. Note that this involves choosing several vertices and colors, but certain pairs of the vertices we choose must induce edges in $E^{\prime}$ and some pairs of vertices and colors must be in $V^{\prime}$. The table also summarizes these requirements.
@@
Table 1. Counting vertices, colors, sharable edges and non-hitting colors.
Each vertex and each color has $O(n)$ choices. The event that any particular pair of vertices induces an edge in $E^{\prime}$ is $p$, which is also the probability of any particular pair $(v,c)$ being in $V^{\prime}$. Thus, the expected number of ways to complete our Type 1 conflict is $O(n^{14+7}p^{24+18})=O(d^{3})$. In a similar approach to that outlined in the proof of Claim 2, McDiarmid’s inequality shows concentration and so w.h.p. each gadget $\Gamma_{1}$ is in $O(d^{3})$ conflicts of Type 1. Notice that we needed to include factors of $p$ in our estimate. Oftentimes, such as in Claim 4, we do not need to be as precise in order to verify the required condition.
Now we claim that each gadget is in $O(d^{3})$ conflicts of Types 2, 3, and 4 (i.e. the rest of the conflicts in $\mathcal{C}_{A}^{(4)}$). These conflicts all involve 4 gadgets such that the sum of the number of vertex choices and color choices is 30. Indeed, for Type 2 and 3 conflicts there are 20 vertices spanned and 10 colors used, while for Type 4 conflicts there are 19 vertices spanned and 11 colors used. Any of these conflicts involves having 32 pairs of vertices in $E^{\prime}$ and 24 vertex-color pairs in $V^{\prime}$. Fixing one of the gadgets gives us 6 vertices, 3 colors, 8 pairs in $E^{\prime}$ and 6 pairs in $V^{\prime}$. Thus the expected number of conflicts of Type 2, 3, or 4 containing a fixed gadget is $O(n^{30-6-3}p^{32+24-8-6})=O(d^{3})$. As before, McDiarmid’s inequality shows concentration.
We move on to $\mathcal{C}_{A}^{(5)}$, meaning Type 5 conflicts. We claim that each gadget is in $O(d^{4})$ conflicts of Type 5. Indeed, such a conflict involves 5 gadgets spanning a total of 24 vertices, using a total of 13 colors, having 40 pairs of vertices in $E^{\prime}$ and 30 vertex-color pairs in $V^{\prime}$. Thus the expected number of conflicts of Type 5 containing a fixed gadget is $O(n^{24+13-6-3}p^{40+30-8-6})=O(d^{4})$.
Finally we consider $\mathcal{C}_{A}^{(6)}$, meaning Type 6 conflicts. These conflicts involve 6 gadgets spanning 29 vertices, using 15 colors, having 48 pairs in $E^{\prime}$ and 36 pairs in $V^{\prime}$. Thus the expected number of conflicts of Type 5 containing a fixed gadget is $O(n^{29+15-6-3}p^{48+36-8-6})=O(d^{5})$.
@@
As we discussed in Section 4.3, we have some leftover uncolored edges after Phase A (this includes both the sharable edges and some more edges). We plan to color them in Phase B by applying Theorem 3 to a new hypergraph $\mathcal{H}_{B}$ and conflict system $\mathcal{C}_{B}$ which together will encode the coloring problem we have for the remaining uncolored edges. However, for the purposes of checking that $\mathcal{H}_{B},\mathcal{C}_{B}$ satisfy the conditions of Theorem 3 we must establish some quasirandom properties of the coloring produced in Phase A. The following lemma summarizes these properties. Recall that $E^{\prime\prime}$ is the set of sharable edges. Let $E^{\prime\prime\prime}$ be the set of non-sharable edges which do not get colored in Phase A. We say the edges in $E^{\prime\prime}\cup E^{\prime\prime\prime}$ are the uncolored edges, and the uncolored degree of a vertex is the number of incident uncolored edges. For $e\in E^{\prime\prime}\cup E^{\prime\prime\prime}$, let $\mathcal{S}_{a,b}(e)$ be the family of sets $S\subseteq V(K_{n})$ with $|S|=a$, having $b$ color repetitions in the coloring from Phase A, and containing an uncolored edge $e^{\prime}$ which is disjoint from $e$ and which belongs to the same set (either $E^{\prime\prime}$ or $E^{\prime\prime\prime}$) as $e$. Roughly speaking, these sets $S$ represent potential danger for Phase B in the sense that if $e$ gets the same color as $e^{\prime}$ in Phase B, then $S$ gains a repeat. We focus on the case where $e$ and $e^{\prime}$ are disjoint since we will be able to do Phase B where every color class is a matching. Also, in Phase B we will use disjoint sets of colors for $E^{\prime\prime}$ and $E^{\prime\prime\prime}$ which is why we restrict our attention to the case when $e$ and $e^{\prime}$ are in the same set.

### v2

@@ 1. Introduction
Gomez-Leos, Heath, Parker, Schweider and Zerbib [17] recently proved that $\frac{6}{7}(n-1)\leq f(n,5,8)\leq n+o(n)$ (the upper bound was also independently proved in [2]). In this paper we give an alternate proof for this lower bound, since we believe it illuminates what must be done to get a matching upper bound. But our main contribution is to obtain a $(5,8)$-coloring using $\frac{6}{7}n+o(n)$ colors. To do this we will actually separately apply certain results of Delcourt and Postle [9] as well as Glock, Joos, Kim, Kühn and Lichev [14]. The results in [9] and [14] are quite similar in the big picture, but there are important technical differences that give each one an advantage over the other in certain instances.
@@ 2. Gadget packing and conflict-free hypergraph matchings
Very recently, Glock, Joos, Kim, Kühn, Lichev and Pikhurko [15] proved Conjecture 1 for the case $\ell=6$ by finding the limit. The proof is significantly more complicated but bears a resemblance to the sketch we provided above for Glock’s proof [13] of the case $\ell=5$. In particular, the lower bound in [15] is obtained by packing an appropriately defined gadget, i.e. a hypergraph which has some sharable pairs. The upper bound follows from establishing a few constraints like (2) and (3) and then using linear programming (and the optimal solution to the linear program tells us what gadget to pack). However, to avoid all possible $(6,4)$-configurations, one must avoid certain configurations consisting of several gadgets intersecting in certain ways. For this, a result like Pippinger and Spencer [22] is not sufficient, and so they use conflict-free hypergraph matchings instead.
@@ 3. Lower bound
Consider a $(5,8)$-coloring of $K_{n}$ using color set $C$. We will use this coloring to define a certain partition of the edges $E(K_{n})$. We start with the partition $\mathcal{P}^{\prime}$ into monochromatic components. In a $(5,8)$-coloring a monochromatic component has at most three edges. For the rest of this proof whenever we say “component” we mean a monochromatic component. It is sufficient to assume that there are no monochromatic triangles. If there were a monochromatic (blue) triangle, each edge incident to this triangle would be an isolated edge in $\mathcal{P}^{\prime}$ and no other edge could be blue. Thus, swapping the colors of an edge of the triangle and an incident edge would yield another $(5,8)$-coloring with the same number of colors without the monochromatic triangle. Therefore, each part of $\mathcal{P}^{\prime}$ is either an isolated edge, 2-path, 3-path, or 3-star. For each part $P\in\mathcal{P}^{\prime}$ consisting of a component in color $c$, we say $P$ hits each pair $(v,c)$ where $v$ is a vertex in the component. Clearly, each pair $(v,c)$ is hit by at most one part.
@@
Now we start to describe the merging. First, for any part $P=\{xy,xv,xu\}\in\mathcal{P}^{\prime}$, i.e. a 3-star component, we see that $\mathcal{P}^{\prime}$ also has singleton parts $\{uv\},\{uy\},\{vy\}$ (if any of these edges were in a larger component we would have a $(5,8)$-violation). See Figure 2(1). We merge $P$ with these singletons forming a new part $Q$ with six edges. Note that two distinct 3-stars can share at most one vertex since otherwise we have a $(5,8)$-violation. Thus each singleton part in $\mathcal{P}^{\prime}$ gets merged with at most one 3-star. In addition to all the 10 pairs hit by the parts we merged to form $Q$, we also say that $Q$ hits $(x,\text{green}),(x,\text{red}),(x,\text{orange})$ (see Figure 2(1)). Any scenario in which a pair is hit twice implies a $(5,8)$-violation. Indeed, in such a scenario, we would be reusing some color used in $\{x,y,u,v\}$ and introducing one additional vertex, resulting in 3 repetitions on 5 vertices. Next, we merge all 3-stars with their respective singletons to form a new partition. We abuse notation and call the new partition $\mathcal{P}^{\prime}$, and we will keep doing so after even more merging until we reach our final partition $\mathcal{P}$.
Next, for any part $P=\{ux,xy,yv\}\in\mathcal{P}^{\prime}$, i.e. a 3-path component, we see that $\mathcal{P}^{\prime}$ also has singleton parts $\{uy\},\{xv\}$ (and also $\{uv\}$, but we don’t need that for this merging). See Figure i.e. 2(2). We merge $P$ with these singletons forming a new part $Q$ with five edges. We say the new part hits the additional pairs $(x,\text{green}),(y,\text{red})$. Next, for any pair of parts $P=\{xu,uv\},P^{\prime}=\{xy,yv\}\in\mathcal{P}^{\prime}$, i.e. a pair of 2-paths sharing endpoints, we will merge $P,P^{\prime}$ and the singleton part $\{xv\}$. See Figure i.e. 2(3). We say the new part hits the additional pairs $(u,\text{green}),(y,\text{green})$. Next, for any pair of parts $P=\{xy,yu\},P^{\prime}=\{xu,uv\}\in\mathcal{P}^{\prime}$ we will merge $P,P^{\prime}$ and the singleton part $\{xv\}$. See Figure i.e. 2(4). We say the new part hits the additional pair $(u,\text{green})$.
@@
For $i=1,\ldots,8$ let $x_{i}$ be the number of parts in $\mathcal{P}$ that look like Figure 2$(i)$. Then summing the edges we have
@@
Let $H_{i}$ be the subset of all pairs $(v,c)$ that are hit by a part of $\mathcal{P}$ of type $(i)$ in Figure 2. Notice that the total number of pairs $(v,c)$ hit is $|\cup_{i=1}^{8}H_{i}|\leq n|C|$. Since each pair $(v,c)$ is hit by at most one part of $\mathcal{P}$, $H_{i}\cap H_{j}=\emptyset$ for each $i\not=j$ and so the total number of $(v,c)$ hit is just the sum of the $|H_{i}|$ for $i=1,\ldots,8$. Further, $|H_{i}|$ can be found by referring to Figure 2: $|H_{1}|=13x_{1},\ldots|H_{8}|=2x_{8}$. Thus,
@@
$=\frac{19}{7}x_{1}+\frac{10}{7}x_{2}+\frac{10}{7}x_{3}+\frac{3}{7}x_{4}+\frac{6}{7}x_{5}+\frac{1}{7}x_{6}+\frac{2}{7}x_{8}\geq 0.$ (5)
@@ 4.1. Intuition
We will need to take care how the gadgets in our packing intersect. We need almost all the edges of $K_{n}$ to be covered by edge-disjoint gadgets, and if we are careless we might end up with two gadgets intersecting as shown in Figure 3(2). In that figure the red and blue edges come from one gadget, and the orange and gray from a second gadget. If the two gadgets intersect as shown we see a $(5,8)$-violation. To avoid this kind of complication we will only allow a pair of gadgets to share a pair of vertices if that pair is a nonedge in both gadgets, as shown in Figure 3(3). In order to ensure we can still cover almost all edges of $K_{n}$ with our gadgets, we will need to make sure that there are not too many of these pairs of vertices that get used as nonedges of gadgets (such pairs will never be covered by our packing). Thus, before we try to get our packing, we will designate a small set of edges to be our shareable edges (see Figure 3(1)). When we pack a gadget, we will make sure the non-edges of the gadget are all shareable, and the edges of the gadget are not shareable. This way we can allow two gadgets to share two vertices as long as this pair of vertices is a shareable edge and therefore a non-edge in each of the gadgets. We will call this gadget in Figure 3(1) $\Gamma$. We will define $\Gamma$ more formally later, but it will be a set of edges colored according to the pattern shown in Figure 3(1) but with the colors possibly permuted, with the dotted edges always being sharable. The formal definition of $\Gamma$ will also involve additional rules about how the colors are chosen, which we start to describe next.
@@ 4.3. Overview of the coloring
Our construction will start by identifying our set of shareable edges, and non-touching colors for each vertex. Then we would like to color almost all the edges (excluding shareable edges and a few other edges) by packing edge-disjoint gadgets. These gadgets must be packed to obey the rules for shareable edges and the colors chosen to obey the rules for non-touching colors. We will use two disjoint sets of colors $C_{A1}$ and $C_{A2}$ for these gadgets, and the colors in $C_{A1}$ and $C_{A2}$ will play different roles. For $c\in C_{A1}$, $c$ will only be allowed to play the role of blue in Figure i.e. 2(7), and so the color class for $c$ will consist of vertex-disjoint 1-paths (i.e. a matching in $K_{n}$). For $c\in C_{A2}$, $c$ will only be allowed to play the role of red or green in Figure i.e. 2(7), so the color class for $c$ will consist of vertex-disjoint 2-paths. In fact, to get our asymptotically optimal bound on the number of colors we will need each color in $C_{A1}$ to induce an almost-perfect matching and each color in $C_{A2}$ to induce a set of 2-paths covering almost all vertices. When we pack the gadgets the colors will also be chosen to obey certain other rules to avoid $(5,8)$-violations. This part of our construction will be called Phase A, and the proof that it works will be an application of Theorem 2. In particular we will define an auxiliary hypergraph $\mathcal{H}_{A}$ and conflict system $\mathcal{C}_{A}$ such that a $\mathcal{C}_{A}$-free matching in $\mathcal{H}_{A}$ corresponds to a partial $(5,8)$-coloring. The advantage of using Theorem 2 here is that this theorem has some extensions which we will use (see Section A). These extensions allow us to prove that the coloring obtained in Phase A has some nice quasirandom properties which will allow us to do the next phase. In Phase B we will color all edges that remain uncolored after Phase A. This will be done by applying Theorem 3 to a new hypergraph $\mathcal{H}_{B}$ and conflict system $\mathcal{C}_{B}$.
@@ 5.1. Setting up our application of Theorem 2
We will apply Theorem 2 with $\varepsilon=\varepsilon_{A}$, and at certain points of our proof we will see inequalities that $\delta,\varepsilon_{A}$ must satisfy. We will say $\delta,\varepsilon_{A}$ are chosen such that
@@
The conflict system $\mathcal{C}_{A}$ for $\mathcal{H}_{A}$ of course has vertex set $E(\mathcal{H}_{A})$, i.e. the set of gadgets. The conflicts (edges) of $\mathcal{C}_{A}$ will be certain sets of gadgets which together would create a $(5,8)$-violation or an alternating 4-cycle (a 4-cycle that is alternately colored using 2 colors). We avoid these alternating 4-cycles since their presence impedes the verification of the conditions of Theorem 2 for the other conflicts.
@@
  • Type 5 conflicts (see Figure 5(5)) correspond to collections of 5 gadgets $\Gamma_{1},\ldots,\Gamma_{5}$ making a $(5,8)$-violation $S$ in the following way. $\Gamma_{1}$ has a triangle in $S$. $\Gamma_{2}$ and $\Gamma_{3}$ each have $c$-colored edge (blue) in $S$, and $\Gamma_{4},\Gamma_{5}$ each have a $c^{\prime}$-colored edge (green) in $S$. The $c$-colored edges are disjoint, as are the $c^{\prime}$-colored edges, and there is no alternating 4-cycle in $S$. In particular the $c$- and $c^{\prime}$-colored edges together make an alternately colored path of length 4 $(s_{1},s_{2},\ldots,s_{5})$ where $s_{1},s_{3},s_{5}$ are all on the triangle from $\Gamma_{1}$. This path could be oriented in a few ways with respect to the orientation of the triangle from $\Gamma_{1}$, and the figure shows one possible orientation.
@@
Indeed, by the previous observation $\Gamma$ cannot have a fourth vertex in $S$. For the 3 vertices $\Gamma$ has in $S$, both $c,c^{\prime}$ are either non-touching for these vertices (as is the case for red with $s$ in Figure 6) or else they are already touching due to $\Gamma$ (as is the case for red with $p,r,q$ and for blue with $p,q,r,s$ in Figure 6). Thus $S$ only has one edge which could be colored either $c$ or $c^{\prime}$.
@@
In light of Proposition 1 and the fact that $C_{A}$ has $\frac{6}{7}n+O(n^{1-\delta})$ colors, our proof of Theorem 5 will be done if we show there is a large enough $\mathcal{C}_{A}$-free matching in $\mathcal{H}_{A}$. To do that, we start checking the conditions of Theorem 2.
@@ 5.2. Checking conditions of Theorem 2
We start verifying the claim with $j=4$. We will check one case in more detail and then provide a general argument for the rest. We will check Type 1 conflicts in detail, i.e. we will check that each gadget is contained in $O(d^{3})$ conflicts of Type 1. Fix a gadget $\Gamma_{1}$. A Type 1 conflict in $\mathcal{C}_{A}^{(4)}$ containing $\Gamma_{1}$ contains 3 other gadgets $\Gamma_{2},\Gamma_{3},\Gamma_{4}$ (see Figure 8(1)). Fixing $\Gamma_{1}$ and assuming we choose all the information for $\Gamma_{2},\Gamma_{3},\Gamma_{4}$ in that order, the choices we must make are summarized in Table 1. Note that this involves choosing several vertices and colors, but certain pairs of the vertices we choose must not induce edges in $E^{\prime}$ and some pairs of vertices and colors must not be in $V^{\prime}$. The table also summarizes these requirements.
@@
Table 1. Counting vertices, colors, sharable edges and non-touching colors.
Each vertex and each color has $O(n)$ choices. The event that any particular pair of vertices induces an edge that is not in $E^{\prime}$ is $p$, which is also the probability of any particular pair $(v,c)$ not being in $V^{\prime}$. Thus, the expected number of ways to complete our Type 1 conflict is $O(n^{14+7}p^{24+18})=O(d^{3})$. In a similar approach to that outlined in the proof of Claim 2, McDiarmid’s inequality shows concentration and so w.h.p. each gadget $\Gamma_{1}$ is in $O(d^{3})$ conflicts of Type 1. Notice that we needed to include factors of $p$ in our estimate. Oftentimes, such as in Claim 4, we do not need to be as precise in order to verify the required condition.
Now we claim that each gadget is in $O(d^{3})$ conflicts of Types 2, 3, and 4 (i.e. the rest of the conflicts in $\mathcal{C}_{A}^{(4)}$). These conflicts all involve 4 gadgets such that the sum of the number of vertex choices and color choices is 30. Indeed, for Type 2 and 3 conflicts there are 20 vertices spanned and 10 colors used, while for Type 4 conflicts there are 19 vertices spanned and 11 colors used. Any of these conflicts involves having 32 pairs of vertices not in $E^{\prime}$ and 24 vertex-color pairs not in $V^{\prime}$. Fixing one of the gadgets gives us 6 vertices, 3 colors, 8 pairs not in $E^{\prime}$ and 6 pairs not in $V^{\prime}$. Thus the expected number of conflicts of Type 2, 3, or 4 containing a fixed gadget is $O(n^{30-6-3}p^{32+24-8-6})=O(d^{3})$. As before, McDiarmid’s inequality shows concentration.
We move on to $\mathcal{C}_{A}^{(5)}$, meaning Type 5 conflicts. We claim that each gadget is in $O(d^{4})$ conflicts of Type 5. Indeed, such a conflict involves 5 gadgets spanning a total of 24 vertices, using a total of 13 colors, having 40 pairs of vertices not in $E^{\prime}$ and 30 vertex-color pairs not in $V^{\prime}$. Thus the expected number of conflicts of Type 5 containing a fixed gadget is $O(n^{24+13-6-3}p^{40+30-8-6})=O(d^{4})$.
Finally we consider $\mathcal{C}_{A}^{(6)}$, meaning Type 6 conflicts. These conflicts involve 6 gadgets spanning 29 vertices, using 15 colors, having 48 pairs not in $E^{\prime}$ and 36 pairs not in $V^{\prime}$. Thus the expected number of conflicts of Type 5 containing a fixed gadget is $O(n^{29+15-6-3}p^{48+36-8-6})=O(d^{5})$.
@@
As we discussed in Section 4.3, we have some leftover uncolored edges after Phase A (this includes both the sharable edges and some more edges). We plan to color them in Phase B by applying Theorem 3 to a new hypergraph $\mathcal{H}_{B}$ and conflict system $\mathcal{C}_{B}$ which together will encode the coloring problem we have for the remaining uncolored edges. However, for the purposes of checking that $\mathcal{H}_{B},\mathcal{C}_{B}$ satisfy the conditions of Theorem 3 we must establish some quasirandom properties of the coloring produced in Phase A. The following lemma summarizes these properties. Recall that $E^{\prime\prime}$ is the set of sharable edges. Let $E^{\prime\prime\prime}$ be the set of non-sharable edges which do not get colored in Phase A. We say the edges in $E^{\prime\prime}\cup E^{\prime\prime\prime}$ are the uncolored edges, and the uncolored degree of a vertex is the number of incident uncolored edges.
Definition 4.
For $e\in E^{\prime\prime}\cup E^{\prime\prime\prime}$, let $\mathcal{S}_{a,b}(e)$ be the family of sets $S\subseteq V(K_{n})$ with $|S|=a$, having $b$ color repetitions in the coloring from Phase A, and containing both $e$ and an uncolored edge $e^{\prime}$ which is disjoint from $e$ and which belongs to the same set (either $E^{\prime\prime}$ or $E^{\prime\prime\prime}$) as $e$.
Roughly speaking, these sets $S$ represent potential danger for Phase B in the sense that if $e$ gets the same color as $e^{\prime}$ in Phase B, then $S$ gains a repeat. We focus on the case where $e$ and $e^{\prime}$ are disjoint since we will be able to do Phase B where every color class is a matching. Also, in Phase B we will use disjoint sets of colors for $E^{\prime\prime}$ and $E^{\prime\prime\prime}$ which is why we restrict our attention to the case when $e$ and $e^{\prime}$ are in the same set. Note that $\mathcal{S}_{4,b}(e)=\emptyset$ for $b\geq 2$. Indeed, the only way for a set $S$ of 4 vertices to get 2 repeats in Phase A is if some gadget has 4 vertices ( like say $\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}$ in Figure 4) in $S$. But in that case, the only two uncolored edges in $S$ share an endpoint, so $S\notin\mathcal{S}_{4,2}(e)$. Also, $\mathcal{S}_{5,b}(e)=\emptyset$ for $b\geq 3$ since Phase A gives a partial $(5,8)$-coloring.